OpenStudy (anonymous):

Write as a sum, difference, or multiple of logarithms: log cubed root a^2B/C

OpenStudy (anonymous):

\[\log \sqrt[3]{\frac{ a^{2} b}{ c}}\]

OpenStudy (bibby):

rewrite the fraction using fractional exponents like so: \[\huge (\frac{ a^2b }{ c })^{\frac{1}{3}}\]